\begin{thebibliography}{105}
\providecommand{\natexlab}[1]{#1}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi:\discretionary{}{}{}#1}\else
  \providecommand{\doi}{doi:\discretionary{}{}{}\begingroup
  \urlstyle{rm}\Url}\fi
\providecommand{\selectlanguage}[1]{\relax}
\providecommand{\bibAnnoteFile}[1]{%
  \IfFileExists{#1}{\begin{quotation}\noindent\textsc{Key:} #1\\
  \textsc{Annotation:}\ \input{#1}\end{quotation}}{}}
\providecommand{\bibAnnote}[2]{%
  \begin{quotation}\noindent\textsc{Key:} #1\\
  \textsc{Annotation:}\ #2\end{quotation}}

\bibitem[{Aguilar et~al.(2014)Aguilar, Santamar{\'\i}a~Bonfil, Froese, and
  Gershenson}]{Aguilar2014The-Past-Presen}
Aguilar, W., Santamar{\'\i}a~Bonfil, G., Froese, T., and Gershenson, C. (2014).
\newblock The past, present, and future of artificial life.
\newblock \emph{Frontiers in Robotics and AI} 1.
\newblock \doi{10.3389/frobt.2014.00008}
\bibAnnoteFile{Aguilar2014The-Past-Presen}

\bibitem[{Amoretti and Gershenson(2016)}]{Amoretti2015Measuring-the-c}
Amoretti, M. and Gershenson, C. (2016).
\newblock Measuring the complexity of adaptive peer-to-peer systems.
\newblock \emph{Peer-to-Peer Networking and Applications} 9, 1031--1046.
\newblock \doi{10.1007/s12083-015-0385-4}
\bibAnnoteFile{Amoretti2015Measuring-the-c}

\bibitem[{Ashby(1956)}]{Ashby1956}
Ashby, W.~R. (1956).
\newblock \emph{An Introduction to Cybernetics} (London: Chapman \& Hall)
\bibAnnoteFile{Ashby1956}

\bibitem[{Atlan and Cohen(1998)}]{AtlanCohen1998}
Atlan, H. and Cohen, I.~R. (1998).
\newblock Immune information, self-organization and meaning.
\newblock \emph{Int. Immunol.} 10, 711---717
\bibAnnoteFile{AtlanCohen1998}

\bibitem[{Batty(2012)}]{Batty:2012Cities}
Batty, M. (2012).
\newblock Building a science of cities.
\newblock \emph{Cities} 29, S9--S16.
\newblock \doi{10.1016/j.cities.2011.11.008}
\bibAnnoteFile{Batty:2012Cities}

\bibitem[{Batty et~al.(2012)Batty, Morphet, Massuci, and Stanilov}]{Batty:2012}
Batty, M., Morphet, R., Massuci, P., and Stanilov, K. (2012).
\newblock Entropy, complexity and spatial information.
\newblock CASA Working Paper 185
\bibAnnoteFile{Batty:2012}

\bibitem[{Bedau and Humphreys(2008)}]{bedau2008emergence}
Bedau, M.~A. and Humphreys, P. (eds.) (2008).
\newblock \emph{Emergence: Contemporary readings in philosophy and science}
  (Cambridge, MA, USA: MIT Press)
\bibAnnoteFile{bedau2008emergence}

\bibitem[{Beer(2000)}]{Beer2000}
Beer, R.~D. (2000).
\newblock Dynamical approaches to cognitive science.
\newblock \emph{Trends in Cognitive Sciences} 4, 91 -- 99.
\newblock \doi{https://doi.org/10.1016/S1364-6613(99)01440-0}
\bibAnnoteFile{Beer2000}

\bibitem[{Bernstein et~al.(2012)Bernstein, Klein, and
  Malone}]{Bernstein2012GlobalBrain}
Bernstein, A., Klein, M., and Malone, T.~W. (2012).
\newblock Programming the global brain.
\newblock \emph{Commun. ACM} 55, 41--43.
\newblock \doi{10.1145/2160718.2160731}
\bibAnnoteFile{Bernstein2012GlobalBrain}

\bibitem[{Bitbol(2012)}]{Bitbol2012}
Bitbol, M. (2012).
\newblock Downward causation without foundations.
\newblock \emph{Synthese} 185, 233--255.
\newblock \doi{10.1007/s11229-010-9723-5}
\bibAnnoteFile{Bitbol2012}

\bibitem[{Bongard and Levin(2021)}]{10.3389/fevo.2021.650726}
Bongard, J. and Levin, M. (2021).
\newblock Living things are not (20th century) machines: Updating mechanism
  metaphors in light of the modern science of machine behavior.
\newblock \emph{Frontiers in Ecology and Evolution} 9, 147.
\newblock \doi{10.3389/fevo.2021.650726}
\bibAnnoteFile{10.3389/fevo.2021.650726}

\bibitem[{B{\"o}rner et~al.(2005)B{\"o}rner, Dall'Asta, Ke, and
  Vespignani}]{borner2005studying}
B{\"o}rner, K., Dall'Asta, L., Ke, W., and Vespignani, A. (2005).
\newblock Studying the emerging global brain: Analyzing and visualizing the
  impact of co-authorship teams.
\newblock \emph{Complexity} 10, 57--67
\bibAnnoteFile{borner2005studying}

\bibitem[{Campbell(1974)}]{Campbell1974}
Campbell, D.~T. (1974).
\newblock `{Downward} causation' in hierarchically organized biological
  systems.
\newblock In \emph{Studies in the Philosophy of Biology}, eds. F.~J. Ayala and
  T.~Dobzhansky (Macmillan). 179--186
\bibAnnoteFile{Campbell1974}

\bibitem[{Chemero(2009)}]{Chemero2009}
Chemero, A. (2009).
\newblock \emph{Radical Embodied Cognitive Science} (Cambridge, MA: The MIT
  Press)
\bibAnnoteFile{Chemero2009}

\bibitem[{Chialvo and Millonas(1995)}]{Chialvo:1995}
Chialvo, D. and Millonas, M. (1995).
\newblock {How swarms build cognitive maps}.
\newblock In \emph{The biology and technology of intelligent autonomous
  agents}, ed. L.~Steels, vol. 144. 439--450
\bibAnnoteFile{Chialvo:1995}

\bibitem[{Clark and Chalmers(1998)}]{clark1998extended}
Clark, A. and Chalmers, D. (1998).
\newblock The extended mind.
\newblock \emph{analysis} 58, 7--19
\bibAnnoteFile{clark1998extended}

\bibitem[{Couzin(2009)}]{Couzin:2009}
Couzin, I.~D. (2009).
\newblock Collective cognition in animal groups.
\newblock \emph{Trends in Cognitive Sciences} 13, 36 -- 43.
\newblock \doi{10.1016/j.tics.2008.10.002}
\bibAnnoteFile{Couzin:2009}

\bibitem[{Cover and Thomas(2006)}]{CoverThomas2006}
Cover, T.~M. and Thomas, J.~A. (2006).
\newblock \emph{Elements of Information Theory} (Wiley-Interscience)
\bibAnnoteFile{CoverThomas2006}

\bibitem[{Davis(2021)}]{10.3389/fcomp.2021.681416}
Davis, M. (2021).
\newblock The brain-as-computer metaphor.
\newblock \emph{Frontiers in Computer Science} 3, 41.
\newblock \doi{10.3389/fcomp.2021.681416}
\bibAnnoteFile{10.3389/fcomp.2021.681416}

\bibitem[{{De Duve}(2003)}]{DeDuve2003}
{De Duve}, C. (2003).
\newblock \emph{Live Evolving: Molecules, Mind, and Meaning} (Oxford University
  Press)
\bibAnnoteFile{DeDuve2003}

\bibitem[{DeCanio and Watkins(1998)}]{DeCanioWatkins1998}
DeCanio, S.~J. and Watkins, W.~E. (1998).
\newblock Information processing and organizational structure.
\newblock \emph{Journal of Economic Behavior and Organization} 36, 275--294
\bibAnnoteFile{DeCanioWatkins1998}

\bibitem[{Dennett(1989)}]{dennett1989intentional}
Dennett, D.~C. (1989).
\newblock \emph{The intentional stance} (MIT press)
\bibAnnoteFile{dennett1989intentional}

\bibitem[{Downing(2015)}]{Downing2015Intelligence-Em}
Downing, K.~L. (2015).
\newblock \emph{Intelligence Emerging: Adaptivity and Search in Evolving Neural
  Systems} (Cambridge, MA, USA: MIT Press)
\bibAnnoteFile{Downing2015Intelligence-Em}

\bibitem[{Dunbar(1993)}]{Dunbar1993}
Dunbar, R. I.~M. (1993).
\newblock Coevolution of neocortical size, group size and language in humans.
\newblock \emph{Behavioral and Brain Sciences} 16, 681--735
\bibAnnoteFile{Dunbar1993}

\bibitem[{Dunbar(2003)}]{Dunbar2003}
Dunbar, R. I.~M. (2003).
\newblock The social brain: mind, language and society in evolutionary
  perspective.
\newblock \emph{Ann. Rev. Anthrop} 32, 163--181
\bibAnnoteFile{Dunbar2003}

\bibitem[{Epstein(2016)}]{Epstein2016}
Epstein, R. (2016).
\newblock The empty brain.
\newblock \emph{Aeon}
\bibAnnoteFile{Epstein2016}

\bibitem[{Equihua et~al.(2020)Equihua, Espinosa~Aldama, Gershenson,
  L{\'o}pez-Corona, Mungu{\'\i}a, P{\'e}rez-Maqueo et~al.}]{10.7717/peerj.8533}
Equihua, M., Espinosa~Aldama, M., Gershenson, C., L{\'o}pez-Corona, O.,
  Mungu{\'\i}a, M., P{\'e}rez-Maqueo, O., et~al. (2020).
\newblock Ecosystem antifragility: beyond integrity and resilience.
\newblock \emph{PeerJ} 8, e8533.
\newblock \doi{10.7717/peerj.8533}
\bibAnnoteFile{10.7717/peerj.8533}

\bibitem[{Escalona-Mor\'{a}n et~al.(2012)Escalona-Mor\'{a}n, Paredes, and
  Cosenza}]{Escalona-Moran:2012}
Escalona-Mor\'{a}n, M., Paredes, G., and Cosenza, M.~G. (2012).
\newblock Complexity, information transfer and collective behavior in chaotic
  dynamical networks.
\newblock \emph{International Journal of Applied Mathematics and Statistics}
  26, 58--66
\bibAnnoteFile{Escalona-Moran:2012}

\bibitem[{Farnsworth et~al.(2013)Farnsworth, Nelson, and
  Gershenson}]{Farnsworth2013Living-is-Infor}
Farnsworth, K.~D., Nelson, J., and Gershenson, C. (2013).
\newblock Living is information processing: From molecules to global systems.
\newblock \emph{Acta Biotheoretica} 61, 203--222.
\newblock \doi{10.1007/s10441-013-9179-3}
\bibAnnoteFile{Farnsworth2013Living-is-Infor}

\bibitem[{Fern{\'a}ndez et~al.(2017)Fern{\'a}ndez, Aguilar,
  Pi{\~n}a-Garc{\'\i}a, and Gershenson}]{FernandezCxLakes}
Fern{\'a}ndez, N., Aguilar, J., Pi{\~n}a-Garc{\'\i}a, C.~A., and Gershenson, C.
  (2017).
\newblock Complexity of lakes in a latitudinal gradient.
\newblock \emph{Ecological Complexity} 31, 1--20.
\newblock \doi{10.1016/j.ecocom.2017.02.002}
\bibAnnoteFile{FernandezCxLakes}

\bibitem[{Fern\'andez et~al.(2014)Fern\'andez, Maldonado, and
  Gershenson}]{Fernandez2013Information-Mea}
Fern\'andez, N., Maldonado, C., and Gershenson, C. (2014).
\newblock Information measures of complexity, emergence, self-organization,
  homeostasis, and autopoiesis.
\newblock In \emph{Guided Self-Organization: Inception}, ed. M.~Prokopenko
  (Berlin Heidelberg: Springer), vol.~9 of \emph{Emergence, Complexity and
  Computation}. 19--51.
\newblock \doi{10.1007/978-3-642-53734-9_2}
\bibAnnoteFile{Fernandez2013Information-Mea}

\bibitem[{Flack(2017)}]{Flack2017}
Flack, J.~C. (2017).
\newblock Coarse-graining as a downward causation mechanism.
\newblock \emph{Philosophical Transactions of the Royal Society A:
  Mathematical, Physical and Engineering Sciences} 375, 20160338.
\newblock \doi{10.1098/rsta.2016.0338}
\bibAnnoteFile{Flack2017}

\bibitem[{Fodor and Pylyshyn(1988)}]{fodor1988connectionism}
Fodor, J.~A. and Pylyshyn, Z.~W. (1988).
\newblock Connectionism and cognitive architecture: A critical analysis.
\newblock \emph{Cognition} 28, 3--71
\bibAnnoteFile{fodor1988connectionism}

\bibitem[{Froese and Stewart(2010)}]{Froese2010}
Froese, T. and Stewart, J. (2010).
\newblock {Life After Ashby: Ultrastability and the Autopoietic Foundations of
  Biological Autonomy}.
\newblock \emph{Cybernetics and Human Knowing} 17, 7--50
\bibAnnoteFile{Froese2010}

\bibitem[{Froese and Ziemke(2009)}]{Froese2009}
Froese, T. and Ziemke, T. (2009).
\newblock Enactive artificial intelligence: Investigating the systemic
  organization of life and mind.
\newblock \emph{Artificial Intelligence} 173, 366--500
\bibAnnoteFile{Froese2009}

\bibitem[{G{\"a}rdenfors(2000)}]{Gardenfors2000}
G{\"a}rdenfors, P. (2000).
\newblock \emph{Conceptual Spaces: The Geometry of Thought}.
\newblock Bradford Books (MIT Press)
\bibAnnoteFile{Gardenfors2000}

\bibitem[{Garnier et~al.(2007)Garnier, Gautrais, and Theraulaz}]{Garnier:2007}
Garnier, S., Gautrais, J., and Theraulaz, G. (2007).
\newblock The biological principles of swarm intelligence.
\newblock \emph{Swarm Intelligence} 1, 3--31.
\newblock \doi{10.1007/s11721-007-0004-y}
\bibAnnoteFile{Garnier:2007}

\bibitem[{Gell-Mann and Lloyd(1996)}]{GellMann:1996}
Gell-Mann, M. and Lloyd, S. (1996).
\newblock Information measures, effective complexity, and total information.
\newblock \emph{Complexity} 2, 44--52.
\newblock \doi{10.1002/(SICI)1099-0526(199609/10)2:1<44::AID-CPLX10>3.0.CO;2-X}
\bibAnnoteFile{GellMann:1996}

\bibitem[{Gershenson(2002)}]{Gershenson2002ua}
Gershenson, C. (2002).
\newblock Contextuality: A philosophical paradigm, with applications to
  philosophy of cognitive science.
\newblock POCS Essay, COGS, University of Sussex
\bibAnnoteFile{Gershenson2002ua}

\bibitem[{Gershenson(2004)}]{Gershenson2004}
Gershenson, C. (2004).
\newblock Cognitive paradigms: Which one is the best?
\newblock \emph{Cognitive Systems Research} 5, 135--156
\bibAnnoteFile{Gershenson2004}

\bibitem[{Gershenson(2007)}]{GershensonDCSOS}
Gershenson, C. (2007).
\newblock \emph{Design and Control of Self-organizing Systems} (Mexico: CopIt
  Arxives).
\newblock Http://tinyurl.com/DCSOS2007
\bibAnnoteFile{GershensonDCSOS}

\bibitem[{Gershenson(2010)}]{Gershenson:2010b}
Gershenson, C. (2010).
\newblock Computing networks: A general framework to contrast neural and swarm
  cognitions.
\newblock \emph{Paladyn, Journal of Behavioral Robotics} 1, 147--153.
\newblock \doi{10.2478/s13230-010-0015-z}
\bibAnnoteFile{Gershenson:2010b}

\bibitem[{Gershenson(2011{\natexlab{a}})}]{Gershenson2011}
Gershenson, C. (2011{\natexlab{a}}).
\newblock \emph{Are Minds Computable?}
\newblock Tech. Rep. 2011.08, Centro de Ciencias de la Complejidad
\bibAnnoteFile{Gershenson2011}

\bibitem[{Gershenson(2011{\natexlab{b}})}]{Gershenson:2010a}
Gershenson, C. (2011{\natexlab{b}}).
\newblock The sigma profile: A formal tool to study organization and its
  evolution at multiple scales.
\newblock \emph{Complexity} 16, 37--44.
\newblock \doi{10.1002/cplx.20350}
\bibAnnoteFile{Gershenson:2010a}

\bibitem[{Gershenson(2012)}]{Gershenson:2007}
Gershenson, C. (2012).
\newblock The world as evolving information.
\newblock In \emph{Unifying Themes in Complex Systems}, eds. A.~Minai,
  D.~Braha, and Y.~{Bar-Yam} (Berlin Heidelberg: Springer), vol. VII. 100--115.
\newblock \doi{10.1007/978-3-642-18003-3_10}
\bibAnnoteFile{Gershenson:2007}

\bibitem[{Gershenson(2013{\natexlab{a}})}]{Gershenson:2011e}
Gershenson, C. (2013{\natexlab{a}}).
\newblock The implications of interactions for science and philosophy.
\newblock \emph{Foundations of Science} 18, 781--790.
\newblock \doi{10.1007/s10699-012-9305-8}
\bibAnnoteFile{Gershenson:2011e}

\bibitem[{Gershenson(2013{\natexlab{b}})}]{Gershenson:2013}
Gershenson, C. (2013{\natexlab{b}}).
\newblock Living in living cities.
\newblock \emph{Artificial Life} 19, 401--420.
\newblock \doi{10.1162/ARTL_a_00112}
\bibAnnoteFile{Gershenson:2013}

\bibitem[{Gershenson(2015)}]{Gershenson2015Requisite-Varie}
Gershenson, C. (2015).
\newblock Requisite variety, autopoiesis, and self-organization.
\newblock \emph{Kybernetes} 44, 866--873
\bibAnnoteFile{Gershenson2015Requisite-Varie}

\bibitem[{Gershenson(2020)}]{Gershenson2018}
Gershenson, C. (2020).
\newblock Information in science and {Buddhist} philosophy: Towards a
  non-materialistic worldview.
\newblock In \emph{Vajrayana Buddhism in Russia: Topical Issues of History and
  Sociocultural Analytics}, eds. A.~M. Alekseyev-Apraksin and V.~M. Dronova
  (Moscow, Russia: Almazny Put). 210--218.
\newblock \doi{10.20944/preprints201812.0042.v1}
\bibAnnoteFile{Gershenson2018}

\bibitem[{Gershenson(2021{\natexlab{a}})}]{Gershenson2021:emergence}
Gershenson, C. (2021{\natexlab{a}}).
\newblock Emergence in artificial life.
\newblock ArXiv:2105.03216
\bibAnnoteFile{Gershenson2021:emergence}

\bibitem[{Gershenson(2021{\natexlab{b}})}]{10.1162/isal_a_00402}
Gershenson, C. (2021{\natexlab{b}}).
\newblock {On the Scales of Selves: Information, Life, and Buddhist
  Philosophy}.
\newblock In \emph{ALIFE 2021: The 2021 Conference on Artificial Life}, eds.
  J.~{\v C}ejkov{\'a}, S.~Holler, L.~Soros, and O.~Witkowski (Prague, Czech
  Republic: MIT Press), 2.
\newblock \doi{10.1162/isal_a_00402}
\bibAnnoteFile{10.1162/isal_a_00402}

\bibitem[{Gershenson et~al.(2020)Gershenson, Trianni, Werfel, and
  Sayama}]{Gershenson2020}
Gershenson, C., Trianni, V., Werfel, J., and Sayama, H. (2020).
\newblock Self-organization and artificial life.
\newblock \emph{Artificial Life} 26, 391--408.
\newblock \doi{10.1162/artl{\_}a{\_}00324}
\bibAnnoteFile{Gershenson2020}

\bibitem[{Gleick(2011)}]{gleick2011information}
Gleick, J. (2011).
\newblock \emph{The information: A history, a theory, a flood} (New York:
  Pantheon)
\bibAnnoteFile{gleick2011information}

\bibitem[{Haken(1988)}]{Haken1988}
Haken, H. (1988).
\newblock \emph{Information and Self-organization: A Macroscopic Approach to
  Complex Systems} (Berlin: Springer-Verlag)
\bibAnnoteFile{Haken1988}

\bibitem[{Haken and Portugali(2015)}]{Haken2015Information-Ada}
Haken, H. and Portugali, J. (2015).
\newblock \emph{Information Adaptation: The Interplay Between Shannon
  Information and Semantic Information in Cognition}, vol. XII of
  \emph{SpringerBriefs in Complexity} (Springer)
\bibAnnoteFile{Haken2015Information-Ada}

\bibitem[{Harvey(2019)}]{doi:10.1177/1059712319856882}
Harvey, I. (2019).
\newblock Neurath's boat and the sally-anne test: Life, cognition, matter and
  stuff.
\newblock \emph{Adaptive Behavior} 0, 1059712319856882.
\newblock \doi{10.1177/1059712319856882}
\bibAnnoteFile{doi:10.1177/1059712319856882}

\bibitem[{Heylighen(1999)}]{Heylighen1999}
Heylighen, F. (1999).
\newblock Collective intelligence and its implementation on the web.
\newblock \emph{Computational and Mathematical Theory of Organizations} 5,
  253--280
\bibAnnoteFile{Heylighen1999}

\bibitem[{Heylighen et~al.(2007)Heylighen, Cilliers, and
  Gershenson}]{HeylighenEtAl2007}
Heylighen, F., Cilliers, P., and Gershenson, C. (2007).
\newblock Complexity and philosophy.
\newblock In \emph{Complexity, Science and Society}, eds. J.~Bogg and R.~Geyer
  (Oxford: Radcliffe Publishing). 117--134
\bibAnnoteFile{HeylighenEtAl2007}

\bibitem[{Hidalgo(2015)}]{Hidalgo2015}
Hidalgo, C.~A. (2015).
\newblock \emph{Why Information Grows: The Evolution of Order, from Atoms to
  Economies} (Basic Books)
\bibAnnoteFile{Hidalgo2015}

\bibitem[{H{\"o}lldobler and Wilson(2008)}]{Holldobler2008}
H{\"o}lldobler, B. and Wilson, E.~O. (2008).
\newblock \emph{The Superorganism: The Beauty, Elegance, and Strangeness of
  Insect Societies} (W. W. Norton \& Company)
\bibAnnoteFile{Holldobler2008}

\bibitem[{Hopfield(1994)}]{Hopfield1994}
Hopfield, J.~J. (1994).
\newblock Physics, computation, and why biology looks so different.
\newblock \emph{Journal of Theoretical Biology} 171, 53--60
\bibAnnoteFile{Hopfield1994}

\bibitem[{Hutchins(1995)}]{Hutchins1995}
Hutchins, E. (1995).
\newblock \emph{Cognition in the Wild} (MIT Press)
\bibAnnoteFile{Hutchins1995}

\bibitem[{Kiverstein and Clark(2009)}]{Kiverstein2009}
Kiverstein, J. and Clark, A. (2009).
\newblock Introduction: Mind embodied, embedded, enacted: One church or many?
\newblock \emph{Topoi} 28, 1--7
\bibAnnoteFile{Kiverstein2009}

\bibitem[{Krakauer et~al.(2020)Krakauer, Bertschinger, Olbrich, Flack, and
  Ay}]{Krakauer:2020}
Krakauer, D., Bertschinger, N., Olbrich, E., Flack, J.~C., and Ay, N. (2020).
\newblock The information theory of individuality.
\newblock \emph{Theory in Biosciences} 139, 209--223.
\newblock \doi{10.1007/s12064-020-00313-7}
\bibAnnoteFile{Krakauer:2020}

\bibitem[{Lehn(1990)}]{Lehn:1990}
Lehn, J.-M. (1990).
\newblock Perspectives in supramolecular chemistry---from molecular recognition
  towards molecular information processing and self-organization.
\newblock \emph{Angewandte Chemie International Edition in English} 29,
  1304--1319.
\newblock \doi{10.1002/anie.199013041}
\bibAnnoteFile{Lehn:1990}

\bibitem[{Levin and Dennett(2020)}]{Levin2020}
Levin, M. and Dennett, D.~C. (2020).
\newblock Cognition all the way down.
\newblock \emph{Aeon}
\bibAnnoteFile{Levin2020}

\bibitem[{Lloyd(2001)}]{lloyd2001measures}
Lloyd, S. (2001).
\newblock Measures of complexity: a non-exhaustive list.
\newblock Department of Mechanical Engineering, Massachusetts Institute of
  Technology
\bibAnnoteFile{lloyd2001measures}

\bibitem[{Malone and Bernstein(2015)}]{malone2015handbook}
Malone, T.~W. and Bernstein, M.~S. (eds.) (2015).
\newblock \emph{Handbook of collective intelligence} (Cambridge, MA, USA: MIT
  Press)
\bibAnnoteFile{malone2015handbook}

\bibitem[{Marshall et~al.(2009)Marshall, Bogacz, Dornhaus, Planqu{\'e}, Kovacs,
  and Franks}]{Marshall:2009}
Marshall, J.~A., Bogacz, R., Dornhaus, A., Planqu{\'e}, R., Kovacs, T., and
  Franks, N.~R. (2009).
\newblock On optimal decision-making in brains and social insect colonies.
\newblock \emph{Journal of the Royal Society Interface} \doi{10.1098/
  rsif.2008.0511}
\bibAnnoteFile{Marshall:2009}

\bibitem[{Martin and Reggia(2010)}]{Martin:2010}
Martin, C. and Reggia, J. (2010).
\newblock Self-assembly of neural networks viewed as swarm intelligence.
\newblock \emph{Swarm Intelligence} 4, 1--36.
\newblock \doi{10.1007/s11721-009-0035-7}
\bibAnnoteFile{Martin:2010}

\bibitem[{Maturana and Varela(1980)}]{Maturana1980}
Maturana, H. and Varela, F. (1980).
\newblock \emph{{Autopoiesis and Cognition: The realization of living}}
  (Dordrecht: Reidel Publishing Company)
\bibAnnoteFile{Maturana1980}

\bibitem[{Mayer-Kress and Barczys(1995)}]{mayer1995global}
Mayer-Kress, G. and Barczys, C. (1995).
\newblock The global brain as an emergent structure from the worldwide
  computing network, and its implications for modeling.
\newblock \emph{The information society} 11, 1--27
\bibAnnoteFile{mayer1995global}

\bibitem[{Michel et~al.(2019)Michel, Beck, Block, Blumenfeld, Brown, Carmel
  et~al.}]{Michel2019}
Michel, M., Beck, D., Block, N., Blumenfeld, H., Brown, R., Carmel, D., et~al.
  (2019).
\newblock Opportunities and challenges for a maturing science of consciousness.
\newblock \emph{Nature Human Behaviour} 3, 104--107.
\newblock \doi{10.1038/s41562-019-0531-8}
\bibAnnoteFile{Michel2019}

\bibitem[{Mitchell(2019)}]{mitchell2019artificial}
Mitchell, M. (2019).
\newblock \emph{Artificial intelligence: A guide for thinking humans} (UK:
  Penguin)
\bibAnnoteFile{mitchell2019artificial}

\bibitem[{Murcio et~al.(2015)Murcio, Morphet, Gershenson, and
  Batty}]{Murcio2015}
Murcio, R., Morphet, R., Gershenson, C., and Batty, M. (2015).
\newblock Urban transfer entropy across scales.
\newblock \emph{PLoS ONE} 10, e0133780.
\newblock \doi{10.1371/journal.pone.0133780}
\bibAnnoteFile{Murcio2015}

\bibitem[{Passino et~al.(2008)Passino, Seeley, and Visscher}]{Passino:2008}
Passino, K.~M., Seeley, T.~D., and Visscher, P.~K. (2008).
\newblock Swarm cognition in honey bees.
\newblock \emph{Behavioral Ecology and Sociobiology} 62, 401--414.
\newblock \doi{10.1007/s00265-007-0468-1}
\bibAnnoteFile{Passino:2008}

\bibitem[{Prokopenko et~al.(2009)Prokopenko, Boschetti, and
  Ryan}]{ProkopenkoEtAl2007}
Prokopenko, M., Boschetti, F., and Ryan, A.~J. (2009).
\newblock An information-theoretic primer on complexity, self-organisation and
  emergence.
\newblock \emph{Complexity} 15, 11--28.
\newblock \doi{10.1002/cplx.20249}
\bibAnnoteFile{ProkopenkoEtAl2007}

\bibitem[{Prokopenko et~al.(2011)Prokopenko, Lizier, Obst, and
  Wang}]{Prokopenko:2011}
Prokopenko, M., Lizier, J.~T., Obst, O., and Wang, X.~R. (2011).
\newblock Relating fisher information to order parameters.
\newblock \emph{Phys. Rev. E} 84, 041116.
\newblock \doi{10.1103/PhysRevE.84.041116}
\bibAnnoteFile{Prokopenko:2011}

\bibitem[{{Ram{\'o}n y Cajal}(1899)}]{ramon1899textura}
{Ram{\'o}n y Cajal}, S. (1899).
\newblock \emph{Textura del sistema nervioso del hombre y de los vertebrados:
  estudios sobre el plan estructural y composici{\'o}n histol{\'o}gica de los
  centros nerviosos adicionados de consideraciones fisiol{\'o}gicas fundadas en
  los nuevos descubrimientos}, vol.~1 (Madrid: Moya)
\bibAnnoteFile{ramon1899textura}

\bibitem[{Res{\'e}ndiz-Benhumea et~al.(2021)Res{\'e}ndiz-Benhumea, Sangati,
  Sangati, Keshmiri, and Froese}]{10.3389/fnbot.2021.634085}
Res{\'e}ndiz-Benhumea, G.~M., Sangati, E., Sangati, F., Keshmiri, S., and
  Froese, T. (2021).
\newblock Shrunken social brains? a minimal model of the role of social
  interaction in neural complexity.
\newblock \emph{Frontiers in Neurorobotics} 15, 72.
\newblock \doi{10.3389/fnbot.2021.634085}
\bibAnnoteFile{10.3389/fnbot.2021.634085}

\bibitem[{Reznikova(2007)}]{Reznikova:2007}
Reznikova, Z. (2007).
\newblock \emph{Animal Intelligence From Individual to Social Cognition}
  (Cambridge University Press)
\bibAnnoteFile{Reznikova:2007}

\bibitem[{Roederer(2005)}]{Roederer2005}
Roederer, J.~G. (2005).
\newblock \emph{Information and its Role in Nature} (Heidelberg:
  Springer-Verlag)
\bibAnnoteFile{Roederer2005}

\bibitem[{Roli et~al.(2018)Roli, Villani, Filisetti, and Serra}]{Roli2018}
Roli, A., Villani, M., Filisetti, A., and Serra, R. (2018).
\newblock Dynamical criticality: Overview and open questions.
\newblock \emph{Journal of Systems Science and Complexity} 31, 647--663.
\newblock \doi{10.1007/s11424-017-6117-5}
\bibAnnoteFile{Roli2018}

\bibitem[{Schaller(1997)}]{591665}
Schaller, R. (1997).
\newblock Moore's law: past, present and future.
\newblock \emph{IEEE Spectrum} 34, 52--59.
\newblock \doi{10.1109/6.591665}
\bibAnnoteFile{591665}

\bibitem[{Scharf(2021)}]{Scharf2021}
Scharf, C. (2021).
\newblock \emph{The Ascent of Information: books, bits, genes, machines, and
  life's unending algorithm} (New York, USA: Riverhead Books)
\bibAnnoteFile{Scharf2021}

\bibitem[{Searle(1980)}]{searle1980minds}
Searle, J.~R. (1980).
\newblock Minds, brains, and programs.
\newblock \emph{Behavioral and brain sciences} 3, 417--424
\bibAnnoteFile{searle1980minds}

\bibitem[{Shalf(2020)}]{doi:10.1098/rsta.2019.0061}
Shalf, J. (2020).
\newblock The future of computing beyond moore's law.
\newblock \emph{Philosophical Transactions of the Royal Society A:
  Mathematical, Physical and Engineering Sciences} 378, 20190061.
\newblock \doi{10.1098/rsta.2019.0061}
\bibAnnoteFile{doi:10.1098/rsta.2019.0061}

\bibitem[{Shannon(1948)}]{Shannon1948}
Shannon, C.~E. (1948).
\newblock A mathematical theory of communication.
\newblock \emph{Bell System Technical Journal} 27, 379--423 and 623--656.
\newblock \doi{10.1002/j.1538-7305.1948.tb01338.x}
\bibAnnoteFile{Shannon1948}

\bibitem[{Smolensky(1988)}]{smolensky1988proper}
Smolensky, P. (1988).
\newblock On the proper treatment of connectionism.
\newblock \emph{Behavioral and Brain Sciences} 11, 1--23
\bibAnnoteFile{smolensky1988proper}

\bibitem[{Sol{\'e} et~al.(2016)Sol{\'e}, Amor, Duran-Nebreda, Conde-Pueyo,
  Carbonell-Ballestero, and Monta{\~n}ez}]{SOLE201647}
Sol{\'e}, R., Amor, D.~R., Duran-Nebreda, S., Conde-Pueyo, N.,
  Carbonell-Ballestero, M., and Monta{\~n}ez, R. (2016).
\newblock Synthetic collective intelligence.
\newblock \emph{Biosystems} 148, 47 -- 61.
\newblock \doi{https://doi.org/10.1016/j.biosystems.2016.01.002}.
\newblock What Synthetic Biology can offer to Artificial Intelligence
\bibAnnoteFile{SOLE201647}

\bibitem[{Steels and Brooks(1995)}]{steels1995artificial}
Steels, L. and Brooks, R. (1995).
\newblock \emph{The Artificial Life Route to Artificial Intelligence: Building
  Embodied, Situated Agents} (New York City, NY, USA: Lawrence Erlbaum
  Associates)
\bibAnnoteFile{steels1995artificial}

\bibitem[{Stewart(1995)}]{Stewart1995}
Stewart, J. (1995).
\newblock Cognition = life : Implications for higher-level cognition.
\newblock \emph{Behavioural processes} 35, 311--326
\bibAnnoteFile{Stewart1995}

\bibitem[{Stewart et~al.(2010)Stewart, Gapenne, and {Di Paolo}}]{Enaction2010}
Stewart, J., Gapenne, O., and {Di Paolo}, E.~A. (eds.) (2010).
\newblock \emph{Enaction: Toward a New Paradigm for Cognitive Science} (MIT
  Press)
\bibAnnoteFile{Enaction2010}

\bibitem[{Trianni and Tuci(2009)}]{Trianni-Tuci:09:ecal}
Trianni, V. and Tuci, E. (2009).
\newblock Swarm cognition and artificial life.
\newblock In \emph{Advances in Artificial Life. Proceedings of the 10th
  European Conference on Artificial Life (ECAL 2009)}
\bibAnnoteFile{Trianni-Tuci:09:ecal}

\bibitem[{Turing(1937)}]{Turing:1937}
Turing, A.~M. (1937).
\newblock On computable numbers, with an application to the entscheidungs
  problem.
\newblock \emph{Proceedings of the London Mathematical Society} s2-42, 230--65.
\newblock \doi{10.1112/plms/s2-42.1.230}
\bibAnnoteFile{Turing:1937}

\bibitem[{Turing(1950)}]{Turing:1950}
Turing, A.~M. (1950).
\newblock Computing machinery and intelligence.
\newblock \emph{Mind} 59, 433--460
\bibAnnoteFile{Turing:1950}

\bibitem[{Varela et~al.(1991)Varela, Thompson, and Rosch}]{Varela1991}
Varela, F.~J., Thompson, E., and Rosch, E. (1991).
\newblock \emph{The Embodied Mind: Cognitive Science and Human Experience}
  (Cambridge, MA: MIT Press)
\bibAnnoteFile{Varela1991}

\bibitem[{von Baeyer(2005)}]{vonBaeyer2004}
von Baeyer, H.~C. (2005).
\newblock \emph{Information: The New Language of Science} (Cambridge, MA:
  Harvard University Press)
\bibAnnoteFile{vonBaeyer2004}

\bibitem[{{von Neumann}(1958)}]{von-Neumann1958}
{von Neumann}, J. (1958).
\newblock \emph{The Computer and the Brain} (New Haven, CT, USA: Yale
  University Press)
\bibAnnoteFile{von-Neumann1958}

\bibitem[{Walker(2014)}]{Walker2014}
Walker, S.~I. (2014).
\newblock Top-down causation and the rise of information in the emergence of
  life.
\newblock \emph{Information} 5, 424--439.
\newblock \doi{10.3390/info5030424}
\bibAnnoteFile{Walker2014}

\bibitem[{Wheeler(1990)}]{Wheeler1990}
Wheeler, J.~A. (1990).
\newblock Information, physics, quantum: the search for links.
\newblock In \emph{Complexity, Entropy, and the Physics of Information}, ed.
  W.~H. Zurek (Reading, MA: Perseus Books), vol. VIII of \emph{Santa Fe
  Institute Studies in the Sciences of Complexity}, chap.~19. 309--336
\bibAnnoteFile{Wheeler1990}

\bibitem[{Wittgenstein(1999)}]{Wittgenstein1999}
Wittgenstein, L. (1999).
\newblock \emph{Philosophical Investigations} (Prentice Hall), 3rd edn.
\bibAnnoteFile{Wittgenstein1999}

\bibitem[{Wolfram(2002)}]{Wolfram:2002}
Wolfram, S. (2002).
\newblock \emph{A New Kind of Science} (Champaign, IL, USA: Wolfram Media)
\bibAnnoteFile{Wolfram:2002}

\bibitem[{Zarkadakis(2015)}]{Zarkadakis2015}
Zarkadakis, G. (2015).
\newblock \emph{In Our Own Image: Savior or Destroyer? The History and Future
  of Artificial Intelligence} (Pegasus Books)
\bibAnnoteFile{Zarkadakis2015}

\bibitem[{Zubillaga et~al.(2014)Zubillaga, Cruz, Aguilar, Zapot\'ecatl,
  Fern\'andez, Aguilar et~al.}]{Zubillaga2014Measuring-the-C}
Zubillaga, D., Cruz, G., Aguilar, L.~D., Zapot\'ecatl, J., Fern\'andez, N.,
  Aguilar, J., et~al. (2014).
\newblock Measuring the complexity of self-organizing traffic lights.
\newblock \emph{Entropy} 16, 2384--2407.
\newblock \doi{10.3390/e16052384}
\bibAnnoteFile{Zubillaga2014Measuring-the-C}

\end{thebibliography}

\documentclass[utf8]{frontiersSCNS} 

\usepackage{url,hyperref,lineno,microtype,subcaption}
\usepackage[onehalfspacing]{setspace}

\linenumbers

\def\keyFont{\fontsize{8}{11}\helveticabold }
\def\firstAuthorLast{Gershenson} 
\def\Authors{Carlos Gershenson\,$^{1,2,3,*}$}
\def\Address{$^{1}$Instituto de Investigaciones en Matemáticas Aplicadas y Sistemas,
Universidad Nacional Autónoma de México, Mexico City, 04510, Mexico \\
$^{2}$Centro de Ciencias de la Complejidad,
Universidad Nacional Autónoma de México, Mexico City, 04510, Mexico \\ 
$^{3}$ Lakeside Labs GmbH, Lakeside Park B04, 9020 Klagenfurt am Wörthersee, Austria}
\def\corrAuthor{ Carlos Gershenson\\
	Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas\\
	Universidad Nacional Autónoma de México\\
	A.P. 20-126, 01000, México, CDMX, México}

\def\corrEmail{cgg@unam.mx}

\begin{document}
\onecolumn
\firstpage{1}

\title[Intelligence as information processing]{Intelligence as information processing: brains, swarms, and computers} 

\author[\firstAuthorLast ]{\Authors} 
\address{} 
\correspondance{} 

\extraAuth{}

\maketitle

\begin{abstract}

\section{}
There is no agreed definition of intelligence, so it is problematic to simply ask whether brains, swarms, computers, or other systems are intelligent or not. To compare the potential intelligence exhibited by different cognitive systems, I use the common approach used by artificial intelligence and artificial life: Instead of studying the substrate of systems, let us focus on their organization. This organization can be measured with information. Thus, I apply an informationist epistemology to describe cognitive systems, including brains and computers. This allows me to frame the usefulness and limitations of the brain-computer analogy in different contexts. I also use this perspective to discuss the evolution and ecology of intelligence.

\tiny
 \keyFont{ \section{Keywords:} mind, cognition, intelligence, information, brain, computer, swarm} 
\end{abstract}

\section{Introduction}

In the 1850s, an English newspaper described the growing global telegraph network as a ``nervous system of the planet'' \citep{gleick2011information}. Notice that this was half a century before Santiago Ram\'on y Cajal~\citeyearpar{ramon1899textura} first published his studies on neurons. Still, metaphors have been used since antiquity to describe and try to understand our bodies and our minds~\citep{Zarkadakis2015,Epstein2016}: humans have been described as made of clay (Middle East) or corn (Americas), with flowing humors, like clockwork automata, similar to industrial factories, etc. The most common metaphor in cognitive sciences has been that of describing brains as computers \citep{von-Neumann1958,10.3389/fcomp.2021.681416}.

Metaphors have been used in a broad range of disciplines. For example, in urbanism, there are arguments in favor of changing the dominant narrative of ``cities as machines'' to ``cities as organisms'' \citep{Batty:2012Cities,Gershenson:2013}.

We can have a plethora of discussions on which metaphors are the best. Still, being pragmatic, we can judge metaphors in terms of their usefulness: if they help us understand phenomena or build systems, then they are valuable. Notice that then, depending on the context, different metaphors can be useful for different purposes \citep{Gershenson2004}. For example, in the 1980s, the debate between symbolists/representationists (brain as processing symbols) \citep{fodor1988connectionism} and connectionists (brain as network of simple units) \citep{smolensky1988proper} did not end with a ``winner'' and a ``loser'', as both metaphors (computational, by the way) are useful in different contexts.

There have been several other metaphors used to describe cognition, minds, and brains, each with their advantages and disadvantages \citep{Varela1991,steels1995artificial,clark1998extended,Beer2000,Gardenfors2000,Garnier:2007,Froese2009,Kiverstein2009,Chemero2009,Enaction2010,Froese2010,Downing2015Intelligence-Em,doi:10.1177/1059712319856882}. It is not my purpose to discuss these here, but to notice that there is a rich variety of flavors when it comes to studying cognition. Nevertheless, all of these metaphors can be described in terms of information processing. Since computation can be understood as the transformation of information \citep{Gershenson:2007}, ``computers'', broadly understood as machines that process information can be a useful metaphor to contain and compare other metaphors. Note that the concept of ``machine'' (and thus computer) could also be updated \citep{10.3389/fevo.2021.650726}.

Formally, computation was defined by Turing~\citeyearpar{Turing:1937}. A computable function is that which can be calculated by a Universal Turing Machine (UTM). Still, there are two main limitations of UTMs related to modeling minds \citep{Gershenson2011}:
\begin{enumerate}
\item \textbf{UTMs are closed}. Once a computation begins, there is no change in the program or data, so adaptation during computation is limited. 
\item \textbf{UTMs compute only once they halt}. In other words, outputs depend on a UTM ``finishing its computation''. Still, minds seem to be more continuous than halting. Then the question arises: what function would a mind be computing?
\end{enumerate}

As many have noted, the continuous nature of cognition seems to be closely related to that of the living \citep{Maturana1980,Hopfield1994,Stewart1995,Walker2014}. We have previously studied the ``living as information processing'' \citep{Farnsworth2013Living-is-Infor}, not only at the organism level, but at all relevant scales. Thus, it is natural to use a similar approach to describe intelligence.

In the next section, I present a general notion of information and its limits to study intelligence. Then, I present the advantages of studying intelligence in terms of information processing. Intelligence is not restricted to brains, and swarms are a classic example of this, which can also be described as information processing systems. Before concluding, I exploit the metaphor of ``intelligence as information processing'' to understand its evolution and ecology.

\section{Information}

Shannon~\citeyearpar{Shannon1948} proposed a measure of information in the context of telecommunications, that is equivalent to Boltzmann-Gibbs entropy. This measure characterizes how much a receiver ``learns'' from incoming symbols (usually bits) of a string, based on the probability distribution of previously known/received symbols: if new bits can be completely determined from the past (as in a string with only one repeating symbol), then they carry zero information (because we know that the new symbols will be the same as previous ones). If previous information is useless to predict the next bit (as in a random coin toss), then the bit will carry maximum information. Elaborating on this, Shannon calculated how much redundancy is required to reliably transmit a message over an unreliable (noisy) channel. Even when Shannon's purpose was very specific, the use of information in various disciplines has exploded in recent decades \citep{Haken1988,Wheeler1990,Lehn:1990,GellMann:1996,AtlanCohen1998,DeCanioWatkins1998,vonBaeyer2004,Roederer2005,CoverThomas2006,ProkopenkoEtAl2007,Prokopenko:2011,Batty:2012,Escalona-Moran:2012,Gershenson:2007,Fernandez2013Information-Mea,Zubillaga2014Measuring-the-C,Hidalgo2015,Haken2015Information-Ada,Murcio2015,Amoretti2015Measuring-the-c,FernandezCxLakes,Roli2018,Krakauer:2020,10.7717/peerj.8533,Gershenson2018,10.1162/isal_a_00402,Scharf2021}.

We can say that electronic computers process information \emph{explicitly}, as we can analyze each change of state and information is encoded in a precise physical location. However, humans and other animals process information \emph{implicitly}. For example, we say we have memories, but these are not physically at a specific location. And it seems unfeasible to represent precisely the how information changes in our brains. Still, we do process information, as we can describe ``inputs'' (perceptions) and ``outputs'' (actions).

Shannon assumed that the \emph{meaning} of a message was agreed previously between emitter and receiver. This was no major problem for telecommunications. However, in other contexts, meaning is not a trivial matter. Following Wittgenstein~\citeyearpar{Wittgenstein1999}, we can say that the meaning of information is given by the \emph{use} agents make of it. This has several implications. One is that we can change meaning without changing information (passive information transformation \citep{Gershenson:2007}). Another is the limits on artificial intelligence \citep{searle1980minds,mitchell2019artificial}, as the \emph{use} of information in artificial systems tends to be predefined. Algorithms can ``recognize'' traffic lights or cats in an image, as they are trained for this specific purpose. But the ``meaning'' for computer programs is predefined, \emph{i.e.} what we want the program to do. The quest for an ``artificial general intelligence'' that would go beyond this limit has produced not much more than speculations.

Even if we could simulate in a digital computer all the neurons, molecules, or even elementary particles from a brain, such a simulation would not yield something akin to a mind. On the one hand, \emph{interactions} generate novel information at multiple scales. On the other hand, as mentioned above, \emph{observers} can give different meanings to the same information. In other words, the same ``brain state'' could refer to different ``mental state'' for different people. In a sense, this is related to the failure of Laplace's daemon: even with full information of the states of the components of a system, prediction is limited because interactions generate novel information \citep{Gershenson:2011e}. And this novel information can determine the future production of information at different scales through upward or downward causation \citep{Campbell1974,Bitbol2012,Flack2017}, so all relevant scales should be considered \citep{Gershenson2021:emergence}.

In spite of all its limitations, the computer metaphor can be useful in a particular way. First, the limits on prediction by interactions are related to \emph{computational irreducibility} \citep{Wolfram:2002}. Second, describing brains and minds in terms of information allows us to avoid dualisms. Thus, it becomes natural to use information processing to describe intelligence and its evolution. Finally, information can contain other metaphors and formalisms, so it can be used to compare them and also to exploit their benefits.

\section{Intelligence}

There are several definitions of intelligence, but not a single one that is agreed upon. We have similar situations with the definitions of life \citep{DeDuve2003,Aguilar2014The-Past-Presen}, consciousness \citep{Michel2019},  complexity \citep{lloyd2001measures,HeylighenEtAl2007}, emergence \citep{bedau2008emergence},
and more. These concepts could be said to be of the type ``I know it when I see it'', to quote Potter Stewart.

Still, having no agreed definition is no motive nor excuse for not studying a phenomenon. Moreover, having different definitions for the same phenomenon can give us broader insights than if we stick to a single, narrow, inflexible definition.

Thus, we could define intelligence as ``the art of getting away with it'' (Arturo Frapp\'e), or ``the ability to hold two opposed ideas in mind at the same time and still retain the ability to function. One should, for example, be able to see that things are hopeless and yet be determined to make them otherwise” (F. Scott Fitzgerald). Alan Turing~\citeyearpar{Turing:1950} proposed his famous test to decide whether a machine was intelligent. Generalizing Turing's test, Mario Lagunez suggested that in order to decide whether a system was intelligent, first, the system has to perform an action. Then, an observer has to \emph{judge} whether the action was intelligent or not, according to some criteria. In this sense, there is no intrinsically intelligent behavior. All actions and decisions are contextual \citep{Gershenson2002ua}. Like with meaning, the same action can be intelligent or not, depending on the context and on the judge and their expectations. 

Generalizing, we can define intelligence in terms of information processing: An agent $a$ can be described as intelligent if it transforms information (individual (internal) or environmental (external)) to increase its ``satisfaction'' $\sigma$.

I have previously defined satisfaction $\sigma \in [0,1]$ as the degree to which the goals of an agent have been fulfilled \citep{GershensonDCSOS,Gershenson:2010a}. Certainly, we still require an observer, since we are the ones who define the goals of an agent, its boundaries, its scale, and thus, its satisfaction. Examples of goals are sustainability, survival, happiness, power, control, and understanding. All of these can be described as \emph{information propagation} \citep{Gershenson:2007}: In this context, an intelligent agent will propagate its own information. 

Brains by themselves cannot propagate. But species of animals with brains tend to propagate. In this context, brains are parts of agents that help process information in order to propagate those agents. From this abstract perspective, we can see that such ability is not restricted to brains \citep{Levin2020}. Thus, there are other mechanisms capable of producing intelligent behavior.

\section{Swarms}

There has been much work related to collective intelligence and cognition \citep{Hutchins1995,Heylighen1999,Reznikova:2007,Couzin:2009,malone2015handbook,SOLE201647}. Interestingly, groups of humans, animals or machines do not have a single brain. Thus information processing is distributed. 

A particular case is that of insect swarms \citep{Chialvo:1995,Garnier:2007,Passino:2008,Marshall:2009,Trianni-Tuci:09:ecal,Martin:2010}, where not only information processing is distributed, but also reproduction and selection occur at the colony level \citep{Holldobler2008}. 

To compare the cognitive architectures of brains and swarms, I previously proposed \emph{computing networks} \citep{Gershenson:2010b}. With this formalism, it can be shown that the differences in substrate do not necessarily imply a theoretical difference in cognitive abilities. Nevertheless, in practice, the speed and scalability of information processing of brains is much superior than that of swarms.

Thus, the brain as computer metaphor is not appropriate for studying collective intelligence in general, nor swarm intelligence in particular. However, the intelligence of brains and swarms can be  described in terms of information processing, as an agent $a$ can be an organism or a colony, with its own satisfaction $\sigma$ defined by an external observer.

Another advantage of studying intelligence as information processing is that we can use the same formalism to study intelligence at multiple scales: cellular, multicellular, collective/social, and cultural. Curiously, at the global scale, the brain metaphor has also been used \citep{mayer1995global,borner2005studying,Bernstein2012GlobalBrain}, although its usefulness remains to be demonstrated.

\section{Evolution and ecology}

If we want to have a better understanding of intelligence, we must study how it came to evolve. Intelligence as information-processing can also be useful in this context, as different substrates and mechanisms can be used to exhibit intelligent behavior. 

What could be the ecological pressures that promote the evolution of intelligence? Since environments and ecosystems can also be described in terms of information, we can say that more \emph{complex} environments will promote --- through natural selection --- more complex organisms and species, which will require a more complex intelligence to process the information of their environment and of other organisms and species they interact with \citep{Gershenson:2007}. In this way, the complexity of ecosystems can also be expected to increase though evolution. 

These ideas generalize Dunbar's \citeyearpar{Dunbar1993,Dunbar2003} ``social brain hypothesis'': larger and more complex social groups put a selective pressure on more complex information processing (measured as the neocortex to bodymass ratio), which gives individuals more cognitive capacities to recognize different individuals, remember who can they trust, multiple levels of intentionality \citep{dennett1989intentional}, and so on. In turn, increased cognitive abilities lead to more complex groups, so this cycle reinforces the selection for more intelligent individuals. 

One can make a similar argument using environments instead of social groups: more complex ecosystems put a selective pressure for more intelligent organisms, social groups, and species; as they require greater information-processing capabilities to survive and exploit their environments. This also creates a feedback, where more complex information processing by organisms, groups, and species produce more complex ecosystems. 

However, individuals can ``offload'' their information processing to their group or environment, leading to a decrease in their individual information processing abilities \citep{10.3389/fnbot.2021.634085}. This is to say that intelligence does not always increase. Although there is a selective pressure for intelligence, its cost imposes limits that depend as well on the usefulness of increased cognitive abilities.

Generalizing, we can say that information evolves to have greater control over its own production \citep{Gershenson:2007}. This leads to more complex information-processing, and thus, we can expect intelligence to increase at multiple scales through evolution, independently on the substrates that actually do the information processing. 

Another way of describing the same: information is transformed by different causes. This generates a \emph{variety} of complexity \citep{Ashby1956,Gershenson2015Requisite-Varie}. More complex information requires more complex agents to propagate it, leading to an increase of complexity and intelligence through evolution.

At different scales, since the Big Bang, we have seen an increase of information processing through evolution. In recent decades, this increase has been supraexponential in computers \citep{591665}. Although there are limitations for sustaining this rate of increase \citep{doi:10.1098/rsta.2019.0061}, we can say that the increase of intelligence is a natural tendency of evolution, be it of brains, swarms, or machines. This will not lead to a ``singularity'', but to an increase of the intelligence and complexity of humans, machines, and the ecosystems we create.

\section{Conclusion}

Brains are not essential for intelligence. Plants, swarms, bacterial colonies, robots, societies, and more exhibit intelligence without brains. An understanding of intelligence (and life \citep{Gershenson2020}) independently of its substrate, in terms of information processing, will be more illuminating that focussing only on the mechanisms used by vertebrates and other animals. In this sense, the metaphor of the brain as a computer, is limited more on the side of the brain than on the side of the computer. Brains do process information to exhibit intelligence, but there are several other mechanisms that also process information to exhibit intelligence. Brains are just a particular case, and we can learn a lot from them, but we will learn more if we do not limit our studies to their particular type of cognition.

\section*{Conflict of Interest Statement}

The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.

\section*{Author Contributions}

C.G. conceived and wrote the paper.


\section*{Funding}
This work was supported by UNAM's PAPIIT IN107919 and IV100120 grants.




\bibliographystyle{frontiersinSCNS_ENG_HUMS} 
\bibliography{carlos,refs}

\end{document}